Solve the following system of equations. If the system cannot be solved, state that it cannot be solved and explain why.
3x-2y=4
2x-5y=21

Answer by Alan3354(69443)

3x-2y=4
2x-5y=21
-----------
Multiply eqn 1 by 2, and eqn 2 by 3
6x -4y = 8
6x -15y = 63
Subtract 2 from 1
0x +11y = -55
y = -5
Sub for y into either eqn
3x -2*(-5) = 4
3x + 10 = 4
3x = -6
x = -2
Both check